2 Cost of Preferred Stock. Micro Spinoffs also has preferred stock outstanding. The stock
pays a dividend of $4 per share, and the stock sells for $40. What is the cost of preferred
stock?
3. Calculating WACC. Suppose Micro Spinoffs’s cost of equity is 12.5 percent. What is its
WACC if equity is 50 percent, preferred stock is 20 percent, and debt is 30 percent of total
capital?
1
Expert's answer
2015-02-26T09:27:23-0500
2. D = $4 per share, P = $40.
Cost of preferred stock = D/P = 4/40 = $0.1.
3. Cost of equity is 12.5 percent, equity is 50 percent, preferred stock is 20 percent, and debt is 30 percent of total capital.
WACC = E/V*Re + D/V*Red*(1 - Tc), where:
Re = cost of equity
Rd = cost of debt
E = market value of the firm's equity
D = market value of the firm's debt
V = E + D
E/V = percentage of financing that is equity
D/V = percentage of financing that is debt
Tc = corporate tax rate
As values E and D are not provided, WACC can't be calculated.